OpenStudy (anonymous):

Solve linear equation 4-7/5x greater than equal to 6

OpenStudy (campbell_st):

ok is the inequation \[4 - \frac{7x}{5} \ge 6\]

OpenStudy (anonymous):

yes thats how its set up

OpenStudy (campbell_st):

ok... subtract 4 from both sides as the 1st step

OpenStudy (anonymous):

ok 4 minus 6=2

OpenStudy (campbell_st):

so you have \[-\frac{7x}{5} \ge 2\] next multiply both sides by 5 to remove the fraction

OpenStudy (anonymous):

i subtract 4 from both sides

OpenStudy (campbell_st):

can you type what you got

OpenStudy (anonymous):

how do i multiply -7/5 by 5

OpenStudy (anonymous):

i have 5 x2=10

OpenStudy (campbell_st):

ok